Your firm is planning a 25 year bond issue. The $1000 par value bonds will pay a semi-annual coupon of $59.80. The investment bankers that you have retained for the bond issue estimate that the bonds will sell for $1,183.12. What is the rate of return that your firm is paying on the bonds?
Answer #1

Given that;
Face value of the bond is $1000
Semiannual coupon payment=$59.8
Present value of the bond is $1183.12
Time period is 25 years. As the coupon payment is made semiannually, the number of periods is 25*2=50
We can determine the rate of return using excel as:

1 Face value 2 Present value 3 Payment 4 Number of periods 5 Semiannual rate of return= 6 Annual rate of return= 1000 -1183.1

1 Face value 1000 2 Present value -1183.12 3 Payment 59.8 4 Number of periods 5 Semiannual rate of return= =RATE(N4,N3, N2,N1

Answer: Hence, the annual rate of return is 9.96%